1.Data are collected on the number of tourists to a city (thousand) and total tourism revenue

($million) for a sample of 20 cities in US. According to the following output, what is the

standard deviation of residuals for this estimated regression equation?

Regression Analysis: Tourism Revenue ($million) versus Tourists (thousand)

The regression equation is: Tourism ($million) = 15.747 + 0.32479 Tourists (thousand)

Predictor Coef SE Coef

Constant 15.747 2.455

Tourists 0.32479 0.07547

S = 1.7386 R-Sq = 63.4%

A) 2.455 B) 0.32479 C) 1.7386 D) 0.634 E) 0.02547

2.Based on the context in Question1, what is the t-statistic to test whether the regression slope is

significant?

Regression Analysis: Tourism Revenue ($million) versus Tourists (thousand)

The regression equation is: Tourism ($million) = 15.747 + 0.32479 Tourists (thousand)

Predictor Coef SE Coef

Constant 15.747 2.455

Tourists (thousand) 0.32479 0.07547

S = 1.7386 R-Sq = 63.4%

A) 0.1323 B) 6.4143 C) 20.8652 D) 4.3036 E) 3.7565
Answer #1

(1) According to the following output, what is the standard deviation of residuals for this estimated regression equation

Answer: (C) 1.7386

(2) 2.Based on the context in Question 1, what is the t-statistic to test whether the regression slope is significant?

t = b1/SE = 0.32479/0.07547 = 4.3036

Answer: (D) 4.3036

[Please give me a Thumbs Up if you are satisfied with my answer. If you are not, please comment on it, so I can edit the answer. Thanks.]